OpenStudy (anonymous):

8x^2=-2x+1 describe the number and type of roots?

OpenStudy (anonymous):

I don't quite know what you mean by describe the number, but if you factor the polynomial, you will get (4x-1)(2x+1), so the roots are x=1/4 and x=-1/2